### v1

# The morphology and interface structure of titanium on graphene

Abstract
Titanium (Ti) is an adhesion and contact metal commonly used in nanoelectronics and two-dimensional (2D) materials research. However, when Ti is deposited on graphene (Gr), we obtain dramatically different film morphology depending on the experimental conditions. Through a combination of transmission electron microscopy, Raman spectroscopy, and ab initio density functional theory calculations, we show that the most critical parameters are the number of Gr layers, the nature of the Gr support, and the deposition temperature. Particularly distinctive is the island morphology and large defect density of Ti on monolayer Gr, compared to bilayer or thicker Gr. We propose that this results from structural and mechanical differences between monolayer and thicker Gr flakes, where monolayer Gr is more flexible, exhibits larger surface roughness and therefore lower Ti diffusivity, and is more easily damaged. Our results highlight the extreme sensitivity of Ti morphology on Gr to processing and substrate conditions, allowing us to propose design rules for controlling Ti-Gr interface properties and morphology and to discuss the implications for other technologically relevant metal deposition processes.
I Introduction
Control over the properties of metal thin films is essential for the functionality of electronic, optical, and magnetic devices. As device miniaturization progresses, the need for atomic-level precision in metal thin film structure becomes increasingly important [1]. Factors such as the nucleation density and growth morphology influence thin-film stress and grain boundary density, and hence the electrical properties of both the metal film and the interface [2, 3, 4].
Figure 1: The effect of Gr layer number on the Ti-Gr interface. (a, b) TEM images of suspended (a) 1L and 2L Gr, (b) 3L Gr. Both images are from the same sample with 1 nm Ti deposited. A step edge separates the 1L and 2L regions in (a). (c) Combined diffraction patterns from regions of suspended 1L, 2L, and 3L Gr. Miller-Bravais indices are indicated. (d) DFT calculations of the atomic interface structure between Gr and the Ti(0001) surface. To highlight the interface structure, the Ti atoms at the interface are shown in green and the rest of the Ti atoms are shown in blue, with carbon atoms shown in dark brown.
On two-dimensional (2D) materials, metal thin film quality is particularly critical for applications where the properties and performance of the system depend on the 2D material/metal interface. Notable examples include contact resistance to 2D materials [5, 6], proximity-induced superconductivity between topological insulators and superconductors [7], and spintronics applications [8]. On van der Waals (vdW) surfaces, the formation of a metal thin film follows steps of initial nucleation, growth of islands, and coalescence [9]. Factors that affect the diffusion distance or nucleation sites may thus impact the properties of the thin film that eventually forms. Various factors determine the metal-2D material interaction, including surface roughness that influences diffusion properties [10, 11], reactions of the metal with the substrate [12] or background gases during deposition [13], contamination from the ambient such as water vapor or hydrocarbons affecting metal morphology and diffusion properties [14], and contamination from preparation of the original surface [15, 16]. Understanding how the extrinsic factors influence metal diffusion, nucleation, and growth is necessary for fabricating well-controlled films with good structural and electronic properties. In graphene (Gr), a material with ultrahigh charge carrier mobility [17, 18], controlling thin film properties and metal-2D interactions is especially critical given the applications in high-frequency electronics [19], spintronic devices [20], wide-band photodetection [21], ballistic devices [22], and hydrogen storage [23].
Of particular interest as a metal thin film is titanium (Ti), which is widely used as an adhesion layer in nanoelectronics due to the strong chemical interactions it can have with substrates, and which shows minimal interdiffusion into noble metals such as gold [24, 25]. Furthermore, Ti islands are of interest for optoelectronic applications, such as the bottom-up fabrication of plasmonic nanoparticle arrays for optical devices [26] and catalysis [27]. However, the exact nature of the interaction between Gr and Ti remains debated. For instance, literature is divided on whether or not defects form in Gr after Ti deposition [28, 29, 30]. In addition, it is possible that Ti and Gr react during deposition, forming TiC [31, 32]. However, another study found that TiC formation is due to a reaction between Ti and background gases or contamination in the vacuum system and not a reaction between the Ti and Gr [33].
To elucidate the fundamental factors determining diffusion, nucleation, and growth of Ti on Gr, and to distinguish between underlying effects such as substrate, strain, and roughness, it is useful to consider metal deposition on suspended Gr, i.e., Gr not in direct contact with a substrate. This approach enables the study of intrinsic metal-Gr interactions without extraneous effects from the substrate. Suspended Gr is itself promising for applications in areas such as filtration [34], DNA sequencing [35], and catalyst support [36, 37, 38]. On suspended Gr, investigations of the early-stage growth of Ti reveal that growth takes place in the Volmer-Weber island growth mode and the initial island layers are strained to accommodate the lattice mismatch between Ti and Gr [39]
Here, we examine the interaction between Ti and Gr systematically, as a function of Gr layer thickness, substrate, and temperature. This allows us to elucidate the influence of these different factors on the Ti-Gr interaction and the resulting Ti morphology and Ti-Gr interface structure. We organize our results as follows: (1) We describe the dependence of Ti island morphology and nucleation characteristics on Gr thickness, through a combination of TEM and ab initio density functional theory (DFT) calculations; (2) We determine the effect of deposition temperature on the Ti morphology; (3) We consider substrate effects by comparing results for suspended Gr and SiN-supported Gr; (4) We investigate defect formation during Ti deposition by Raman spectroscopy. (5) We summarize the effects of Gr thickness, temperature, substrate on Ti morphology and defect formation.
II Graphene thickness dependence on Ti-island morphology and nucleation characteristics
Figure 1(a, b) shows the results of depositing, by evaporation, a 1 nm thick Ti layer at room temperature onto 1L, 2L and 3L suspended Gr. Optical microscopy images of the Gr before and after transfer to the TEM sample carrier, shown in Fig. S1, allow us to correlate the location of images recorded in TEM with the thickness of the Gr at that point. We find that the Ti forms irregularly shaped islands on Gr, but in spite of their irregular appearance these islands are aligned epitaxially. An irregular island shape is expected given the low homologous temperature during these depositions [2],  0.15 $T_{\mathrm{M}}$, where $T_{\mathrm{M}}$=1941 K is the melting temperature of Ti.
The uniformity of the Ti film suggests that the growth is not dominated by a low density of heterogeneous nucleation sites such as surface contaminants. In general, to investigate the intrinsic interaction between metals and 2D materials, it is crucial that the 2D material is free from adsorbed contaminants, particularly water and polymer residues from the transfer process. The samples shown in Fig. 1 were prepared using a transfer process combined with sample annealing and metal deposition, the latter two conducted in ultra-high vacuum (UHV) conditions, to create the suspended and supported Gr samples on which we deposit Ti. This process results in the near-complete elimination [10, 40] of contamination on suspended Gr. Even so, it is important to minimize the time between Gr annealing and Ti deposition, because the Gr surface can re-contaminate even in UHV conditions [10, 41]. Further details of sample preparation and Ti deposition are described in Methods.
We now discuss the epitaxial arrangement displayed by the Ti on Gr. The Ti $\langle 2\bar{1}\bar{1}0\rangle$ directions are parallel to the Gr $\langle 10\bar{1}0\rangle$ directions as shown in the diffraction patterns in Fig. 1(c), and the Gr [1000] surface normal is parallel to the Ti [1000] surface normal.
We used a selected area aperture with a diameter of 520 nm. Therefore, each diffraction pattern include signals from several hundred Ti islands. The diffraction patterns show that each separate Ti island grows in identical alignment with the Gr. This crystal alignment between Ti and Gr is consistent with predictions from ab initio calculations that we performed (see Methods), as well as with results from literature [29, 39, 40, 42]. Figure 1(d) shows the expected atomic interface structure between Ti and Gr.
The nucleation density of Ti islands determines parameters such as density of grain boundaries and film stress and has a direct impact on the properties of the resulting thin films [1, 2, 3, 4]. For the Ti islands on suspended Gr in Fig. 1, we find that changes in nucleation density and island size are visible depending on the Gr layer number. Each area shown in Fig. 1(a, b) received the same Ti flux, yet the island size increases and nucleation density decreases with increasing Gr layer number. Figure 2(a) shows this quantitatively in a plot of the Ti-island nucleation density for 1L-3L Gr following the deposition of 1 nm Ti at room temperature, while Fig. 2(b) shows histograms of the island areas. Figure S2 and S3 show the details of the images from which we derived this data and the method used to determine the individual grain boundaries. Details of these measurements are given in Methods.
Figure 2: Ti nucleation characteristics on suspended Gr. (a) Nucleation density measured from the sample shown in Fig. 1. (b) Histograms of Ti island areas for 1L, 2L, and 3L Gr. n is the number of measured islands. (c) Top and side view, respectively, of the most stable adsorption site, the hollow (H) site. Also indicated are the top (T) and bridge (B) sites.
Gr thickness (L) Total adsorption energy (eV) Short-range adsorption energy (eV) vdW adsorption energy (eV) Interface energy (eV/Å2)
1 -2.21 -1.89 -0.32 -0.148
2 -2.33 -1.88 -0.45 -0.161
3 -2.36 -1.91 -0.45 -0.163
Table 1: Results from DFT calculations. The total, short-range, and vdW adsorption energies are for a single Ti atom on the hollow site. The interface energy is calculated for a Ti and Gr slab as outlined in Methods.
We use this data in a mean-field diffusion theory to estimate the diffusion constant of Ti on suspended Gr. Assuming a constant flux of Ti from the evaporator to all regions of the sample, the nucleation density, N, can be expressed as [43]
$N\propto\left(\frac{1}{D}\right)^{\frac{i}{i+1}}$ (1)
where $D$ is the diffusion constant and $i$ is the critical nucleus size. We use values $N_{\mathrm{1L}}$ = 960 µm-2, $N_{\mathrm{2L}}$ = 590 µm-2, and $N_{\mathrm{3L}}$ = 520 µm-2 (Fig. 2). The critical nucleus size of metals can generally be considered small [43, 44], so we assume that $i$=1 or 2. These values lead to results $D_{1}/D_{2}$ = 0.18 and 0.33, respectively, and $D_{1}/D_{3}$ = 0.12 and 0.25, for $i$=1 and 2, respectively. Thus, the measurement shows that the diffusion coefficient of Ti on suspended Gr is 3-5 times larger for 2L Gr compared to 1L Gr, and 4-8 times larger for 3L Gr compared to 1L Gr.
Such layer-dependent metal diffusion properties have previously been observed for gold on Gr [10, 45, 46, 47, 48]. The Gr layer dependence for Au diffusion was attributed to variations in surface roughness with layer number [10, 11]. The underlying mechanism involves the diffusion coefficient being reduced or increased by compressive or tensile strain, respectively [49, 50]. As a result, regions with compressive strain act as effective nucleation sites due to lower metal diffusivity in those areas. Suspended 1L Gr exhibits a root mean square roughness of about 0.1 nm [51, 52, 53], which decreases as the layer thickness increases [54]. Extending these ideas from gold to Ti, the thickness dependence on Ti island size in suspended Gr can be understood consistently in terms of thickness-dependent surface roughness.
To aid in understanding the differences in the Ti nucleation density and interaction as a function of Gr layer number, we performed ab initio DFT calculations. As we show below, these calculations predict an increasing Ti-Gr interaction with increasing Gr thickness. Since experimental results in Fig. 1 and 2 demonstrate the opposite trend, that the Ti-Gr interaction decreases with increasing Gr thickness, our DFT results in fact suggest that extrinsic factors that are not considered in the DFT calculations dominate the layer-dependent Ti-Gr interaction.
Our calculations first consider the single atom adhesion energy of Ti, which was calculated using $E_{\mathrm{ads}}=E_{\mathrm{(Gr/Ti)}}-E_{\mathrm{Gr}}-E_{\mathrm{Ti}}$, where $E_{\mathrm{(Gr/Ti)}}$, $E_{\mathrm{Gr}}$, and $E_{\mathrm{Ti}}$ are the energies of the 1L-3L Gr with adsorbed Ti, pristine Gr, and isolated Ti atoms with a spin configuration consistent with Hund’s rules, respectively. Given the symmetry of the Gr lattice, a Ti adatom can adopt three different adsorption configurations on the top, bridge, and hollow sites, as shown in Fig. 2(c). Our calculations show that, irrespective of the number of Gr layers, the hollow site is the most stable adsorption site for Ti with $E_{\mathrm{ads}}$ for 1L-3L given in Table 1. Adsorption energies for the other sites are given in Table S1. The values are consistent with previous calculations on 1L Gr [55] and are also consistent with adsorption trends of metals on other 2D surfaces such as MoS2 [56, 57].
Table 1 shows that the Ti adsorption energy decreases (i.e. stronger interaction) by $\sim$0.1 eV/Ti atom as the Gr thickness increases from 1L to 2L, and then decreases further by $\sim$0.03 eV/Ti atom from 2L to 3L. Examining the contributions of the short-range and long-range (vdW) adsorption energy shows that the slight increase in the total adsorption strength as Gr thickness increases originates from increased vdW interactions. The short-range interactions, as captured by the energy without the vdW corrections, show the opposite trend.
We next consider the interface energy between Gr and Ti. The interface energy was found by minimizing the misfit strain between Ti (0001) and Gr, with structures as shown in Fig. 1(d). The results for the interface energies are given in Table 1. As is the case for single adsorbed Ti atoms, we find that the interface energy decreases slightly with Gr thickness due to vdW interactions, and hence predicts a stronger Ti-Gr interaction with increasing Gr thickness.
Our overall conclusion from the ab initio calculations for both single Ti atoms and slabs of Ti is that the intrinsic interactions between Gr and Ti are not very sensitive to the Gr thickness, and that a slightly stronger interaction is expected when the Gr thickness increases. This is not consistent with the experimental results in Fig. 1, where weaker interaction, i.e. larger diffusivity was seen with increasing Gr layer number. Therefore, we posit that the differences in nucleation density for 1L-3L suspended Gr that are seen experimentally are due to an extrinsic factor, with surface roughness being a plausible origin.
In summary, we find that increasing Gr thickness leads to higher Ti diffusivity on the Gr surface, resulting in epitaxial islands with a lower nucleation density and hence larger size. We suggest, based on the comparison with DFT calculations, that the observed Gr thickness dependence is due to surface roughness which decreases with increasing Gr thickness, analogous to previous observations for Au on Gr [10]. We summarize our findings in Table 2 along with other data that will be discussed next.
III Effect of deposition temperature on Ti morphology
As a means of Ti morphology control, we now discuss samples where Ti has been deposited on Gr at elevated temperatures. Figure 3(a) shows a TEM image of 1L and 2L Gr after 1 nm Ti deposition while heating to 400 °$\degree$°C. Additional images of 1L, 2L, and 3L Gr with Ti deposited at room temperature, 300 °$\degree$°C, and 400 °$\degree$°C are provided in Fig. S4 and S5.
There is a marked difference in the Ti morphology between 1L and 2L or 3L Gr after a heated Ti deposition. The difference is more evident than that seen for room temperature deposition. On 2L and 3L Gr heated deposition produces faceted islands that approach the Winterbottom equilibrium shape of hexagonally close-packed Ti (inset, Fig. 3(a); see Methods for calculation details) [58]. It is typical for metallic islands to exhibit dendritic growth at lower temperatures and compact, faceted growth at higher temperatures, due to increased adatom diffusion along island edges as well as diminished kinetic limitations [1]. The diffraction pattern for 2L Gr (Fig. 3(b)) shows that the faceted Ti islands are epitaxially aligned with the Gr. Furthermore, we estimate that 15-20% of the islands are rotated 30°$\degree$ with respect to the expected epitaxial alignment. Hence, the effect of increasing the temperature is generally to form compact and faceted islands. This is similar to literature observations of Nb, Au, and Ti on few-layer Gr [40], corresponding to a reduction of the influence of kinetic factors at elevated temperatures.
Figure 3: Layer and substrate dependent Ti morphology after depositing Ti at 400 °$\degree$°C. (a) TEM image of a region with 1L and 2L Gr after Ti deposition at 400 °$\degree$°C. The inset shows the calculated equilibrium Winterbottom shape of Ti which is bounded by the top (0001) surface (red) and the $\{11\bar{2}1\}$ side facets (gold). (b) Diffraction patterns from the 1L and 2L region. Gr and Ti reflections are indexed with black and green font, respectively. The black and green arrow point to a Gr $\{2\bar{1}\bar{1}0\}$ and a second order Ti $\{10\bar{1}0\}$ spot.
On 1L Gr, however, we find that heated deposition produces irregularly shaped islands. This can be understood if reactions with Gr hinder the free movement of the incoming Ti atoms. We have in several instances observed visible damage to the 1L Gr after heated Ti deposition, such as holes or rolled over Gr (see top right corner of the image in Fig. 3(a)), indicating damage to 1L Gr during heated deposition. This is also consistent with Raman spectroscopy data, presented below, showing more defect formation and/or stronger Ti-Gr interaction on 1L Gr. Thus, these results indicate that the suspended 1L Gr becomes damaged when depositing Ti while heating the Gr to 300-400 °$\degree$°C.
Figure 4: The effect of substrate on Ti morphology. (a, b) Depositions at room temperature. AFM scans of 1.1 nm Ti deposited onto regions with (a) suspended, (b) an area containing both suspended and SiN-supported 1L Gr as indicated with arrows. (c, d) Depositions at 300 °$\degree$°C. DF-STEM images of (c) 1L Gr and (d) 2L Gr.
The diffraction pattern for 1L Gr (Fig. 3(b)) also indicates defects in the Gr. There are additional Gr reflections at the positions corresponding to the Gr $\{10\bar{1}0\}$ d-spacing, which indicate the formation of polycrystalline Gr. In reaching this conclusion, we also considered the possibility of a chemical reaction which forms a material with atomic spacing similar to that of the Gr $\{10\bar{1}0\}$ reflections. The most likely candidates would be TiC or TiO2, discussed in Table S2 and its caption. However, comparing our diffraction patterns with expected diffraction reflections from these compounds does not yield a close fit. Other possibilities could include ternary compounds of C, Ti, and O, or the formation of a non-equilibrium compound. However, DF-TEM imaging of 1L Gr shows that the reflections from the diffuse Gr ring do not correlate well with the locations of the Ti islands (Fig. S6). Overall, therefore, we believe the additional reflections at the Gr $\{10\bar{1}0\}$ are most likely due to damage to the Gr and the formation of polycrystalline Gr. An additional observation is that the Ti $\{10\bar{1}0\}$ and $\{2\bar{1}\bar{1}0\}$ spots are more diffuse on suspended 1L Gr compared to 2L Gr, indicating a larger variation in crystal orientations. There is also a second set of faint spots by the Gr $\{2\bar{1}\bar{1}0\}$ spots at a slightly smaller radius (marked with a green arrow, Fig. 3(b)). These match precisely with the second order Ti $\{10\bar{1}0\}$ reflections.
In summary, we find that the result of high temperature deposition depends strongly on the Gr thickness. For 2L and 3L Gr, depositing Ti at 300 or 400 °$\degree$°C results in the formation of faceted and compact islands with morphologies close to the Winterbottom shape, as expected when thermodynamic factors determine island shape. On 1L Gr, depositions at 300 or 400 °$\degree$°C lead to irregularly shaped islands which can be explained by strong interactions with the Gr and defect formation.
IV Substrate dependence of Ti morphology and nucleation characteristics
In most technologically relevant applications, the Gr is supported on a substrate rather than used as a freestanding film. We showed previously that Gr roughness affects Ti nucleation, and it is known that substrates such as SiO2 or SiN increase the surface roughness of Gr [59]. We therefore now discuss Ti morphology on SiN-supported Gr, comparing with freestanding Gr to investigate substrate effects. Figure 4(a, b) compares atomic force microscopy (AFM) images of SiN-supported and suspended 1L Gr with Ti deposited at room temperature. SiN-supported 1L Gr shows a larger nucleation density and smaller island size compared to suspended 1L Gr, to such an extent that identifying individual islands becomes challenging.
We again interpret these results based on differences in surface roughness. In literature for Gr supported on SiO2 and SiN, the root mean square roughness has been measured to be 0.19-0.35 nm, attributed to roughness imposed from these uneven surfaces [59, 60, 61, 62]. This is larger than the intrinsic roughness experienced by suspended 1L Gr, which in literature has been measured to be about 0.1 nm [51, 52, 53]. Therefore, we would expect lower Ti diffusivity on Gr supported by a substrate, which in our case results in a very large nucleation density.
We then compare the Ti island morphology between suspended and SiN-supported Gr after deposition at 300 °$\degree$°C. Figure 4(c, d) shows dark field-STEM images of regions containing both suspended and supported 1L and 2L Gr, respectively. Suspended 1L Gr displays irregularly shaped islands as expected. Interestingly, the SiN-supported 1L Gr displays compact islands which suggests less defect formation in the Gr and a weaker Ti-Gr interaction in these regions.
On 2L Gr, the Ti nucleation density is larger on SiN-supported Gr compared to suspended Gr. This is again indicative of smaller surface roughness on the suspended region, and similar to results obtained in a different study with Au deposition on Gr [10]. We rule out temperature differences between suspended and supported 1L Gr as the reason for the observed differences in Ti island morphology. This is because the whole TEM sample carrier is heated resistively during heated depositions, and we would expect a uniform temperature in the small area where the Gr is located.
Hence, we conclude that the presence of a substrate increases surface roughness, resulting in a larger nucleation density of islands. This is seen for both 1L Gr at room temperature and for 2L Gr at elevated temperatures. Depositions on SiN-supported 1L Gr at elevated temperature result in the formation of compact Ti islands, indicating that the presence of a substrate leads to less defect formation compared to depositions on 1L suspended Gr at elevated temperature where irregularly shaped islands form.
V Defect formation in Gr during Ti deposition
So far, we have discussed nucleation density and Ti island morphology, which are macroscopic phenomena. Now we discuss a key nanoscale structural feature, defects in Gr, which are typically studied by Raman spectroscopy.
Figure 5: Raman spectroscopy. (a) After Ti deposition on suspended Gr without intentional heating. (b) After Ti deposition on suspended Gr heated to 400 °$\degree$°C.
To evaluate defect formation in the Gr during Ti evaporation we obtained Raman spectroscopy data from samples with 1L, 2L, and 3L suspended Gr after Ti deposition at room temperature (Fig. 5(a)). Raman spectra of Gr show characteristic G and 2D peaks at $\sim$1580 cm-1 and 2690 cm-1, respectively. Literature indicates that the D-peak at $\sim$1350 cm-1 appears only in the presence of certain defects. The D-peak is associated with vacancy-type and sp3-type defects such as TiC bonds [63], while charged impurities, intercalants and strain do not produce a D-peak [64]. In our case, suspended 1L Gr with Ti deposited at room temperature exhibits a D-peak, while 2L and 3L Gr do not show a noticeable D-peak. This suggests that defect formation, such as vacancy-type and sp3-type defects, is strongly reduced or even eliminated for 2L and 3L Gr. Similarly, Ti deposition at 400 °$\degree$°C leads to a D-peak on 1L Gr, which is relatively larger than that from 1L Gr room temperature deposition.
The defect concentration is quantified using the intensity ratio between the D- and the G-peak, $I_{\mathrm{D/G}}$ [65]. For 1L Gr at room temperature ($I_{\mathrm{D/G}}$=1.3), the defect density is 1.4$\cdot$104 µm-2. With a Ti nucleation density of 960 µm-2 on 1L suspended Gr at room temperature this suggests that each Ti island contains $\sim$15 point defects in the underlying 1L Gr support. For 1L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}$=2.9), the defect density increases to $5.4\cdot 10^{4}$ µm-2. For 2L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}\sim$0.1), we estimate an upper bound of $<1.6\cdot 10^{3}$ µm-2 defects. This upper bound is based on measurements in Ref. [65], which considers a $I_{\mathrm{D/G}}$-ratio down to 0.2, corresponding to this defect density. These results indicate a strong Gr thickness and temperature dependence on the defect formation in Gr after Ti deposition.
If we compare these Raman spectroscopy data with the structural characterization from Fig. 1 and 3, we obtain a more complete picture of the defect structure of the samples. For room temperature deposition, 1L Gr exhibits many more defects in Raman spectroscopy than 2L or 3L Gr, although structural changes (nucleation density and island size) change more gradually with Gr layer number. In general, high-temperature depositions show more defects compared to room temperature depositions. For high temperature depositions, 1L Gr shows at least an order of magnitude higher defect density than 2L or 3L Gr. For the high temperature case, the structural changes are also dramatic: 1L Gr displays irregularly shaped Ti islands and visible holes, whereas 2L and 3L Gr show compact and faceted islands. Our previous analysis shows that increasing surface roughness results in a stronger Ti-Gr interaction. Since surface roughness of free-standing Gr decreases with increasing thickness [54], our Raman and structural TEM characterization align well with previous studies linking the reactivity of Gr to its surface roughness [66, 67, 68].
VI Conclusions
To conclude, we report a complicated dependence of the Ti-Gr interface morphology and defect formation on layer thickness, substrate, and temperature. We summarize our findings in Table 2. In general, Ti nucleation density depends on the Gr thickness. The island shape and degree of faceting depend on the temperature, in a way expected from kinetic considerations. We found that the Gr thickness dependence on Ti diffusivity and defect formation as measured by Raman spectroscopy depends on surface roughness, where increased surface roughness leads to a stronger Ti-Gr interaction. This was supported by DFT simulations which do not predict a strong Gr thickness dependence on the interaction with Ti for pristine Gr, suggesting that extrinsic factors such as roughness are important.
Parameter Effect of Ti morphology Effect on Gr damage
Gr layer number Increasing Gr thickness leads to larger Ti diffusivity on the Gr surface, resulting in lower nucleation density and larger Ti islands (Fig. 1 & 2). Damage (as identified through Raman spectroscopy) decreases with Gr thickness (Fig. 5).
Temperature Increasing the temperature diminishes kinetic factors and increases thermodynamic factors in the growth of Ti islands, resulting in faceted islands with a shape predicted by the Winterbottom reconstruction (Fig. 3). Increasing the temperature results in more damage to the Gr (Fig. 5).
Substrate A substrate like SiN reduces Ti diffusivity due to increased surface roughness resulting in a continuous film with small grain sizes when depositing Ti at room temperature. At elevated temperature we also observe an increased nucleation density on supported Gr compared to free-standing Gr (Fig. 4). Faceted islands on SiN-supported 1L Gr (as opposed to the irregular islands observed on suspended 1L Gr) after Ti deposition at elevated temperatures suggest less defect formation when a substrate is present (Fig. 4).
Table 2: Summary of experimental results for the dependence of Ti deposition on growth and substrate parameters.
Results for 1L Gr stand out from the results of 2L or 3L Gr. 1L Gr displays at least an order magnitude larger defect density and completely different Ti island morphologies as compared to thicker Gr flakes. We believe these differences arise due to structural and mechanical differences between 1L and thicker Gr flakes, where 1L Gr is more flexible, exhibits larger surface roughness and, therefore, displays lower Ti diffusivity, and is more easily damaged.
These results highlight the extreme sensitivity of processing conditions on Ti deposition on Gr and emphasize the importance of considering factors such as the Gr thickness, the substrate on which Gr is placed, surface roughness, and strain for applications where the microscopic nature of the metal-Gr interface is critical to performance. This is in addition to the usual factors considered in metal nucleation and growth, such as deposition rate and substrate temperature.
The summary in Table 2 offers guidelines for creating Ti-Gr interfaces with ideal properties. For large single crystals with minimized defect formation, it would be interesting to investigate Ti growth on Gr supported by hexagonal boron nitride or another thicker 2D material, as this should suppress damage formation, allowing the growth of facetted islands, while also providing a flat substrate minimizing nucleation density and creating large islands. It may be further possible to grow large, faceted islands by increasing the temperature. Also, interesting would be extending this work to other 2D materials like transition metal dichalcogenides (TMDs), where the Ti-TMD interface is heavily utilized in nanoelectronics research and control of the reactions at the interface are desired [12, 13]. Future studies should also examine the properties of Ti as an adhesion layer to other metals, and how the properties of other metals are affected by the Ti layer. It may be possible to grow a metallic overlayer which is epitaxial with respect to the epitaxial Ti film, and therefore create epitaxial multilayers on Gr. Furthermore, the faceted Ti islands obtained by evaporating Ti at elevated temperatures may be useful for optoelectronic applications and may serve as seeds for other metals on Gr to create complicated multilayer island structures.
VII Methods
VII.1 Sample Fabrication
Gr was obtained by exfoliating natural graphite (NaturGrafit GmbH, Germany) onto oxygen plasma treated substrates of 90 nm thick SiO2 on Si using 3M Magic Scotch tape. We identified suitable flakes by their contrast using optical microscopy as shown in Fig. S1. We then transferred crystals to either homemade TEM compatible sample carriers or ”location-tagged TEM grids” purchased from Norcada, Inc., Canada. The homemade grids consist of 300 nm thick SiN membrane windows with multiple $\sim$3 µm diameter holes etched in the window. The Norcada TEM grids have a 200 nm thick SiN membrane with several 2 µm holes. The Gr is transferred using wedging transfer, based on cellulose acetate butyrate as polymer handle for the transfer, which we have previously shown produces ultra-clean suspended Gr after thermal annealing in vacuum [10, 51]. After the transfer we bake the TEM sample carriers to 140 °$\degree$°C to improve adhesion. Then we dissolve the polymer in acetone and rinse the TEM sample carriers in isopropanol. Finally, we dry the sample with critical point drying to avoid damage to the suspended Gr sample.
VII.2 Ti deposition
For the suspended Gr, to remove polymeric residue, we anneal the sample overnight at 500 °$\degree$°C prior to evaporation. The annealing takes place in an ultrahigh vacuum chamber that also contains the Ti evaporator. We deposit Ti using an electron-beam evaporator at a typical rate of around 0.3 Å/min and a thickness of approximately 1 nm.
VII.3 Electron microscopy
TEM images and diffraction patterns were obtained from a JEOL 2010 TEM operated at 200 kV. A selected area aperture with a diameter of 520 nm was used for acquiring diffraction patterns. DF-STEM and TEM images were obtained from a Hitachi HF5000 environmental TEM operated at 200 kV.
VII.4 Data analysis
Ti island nucleation density was estimated by delineating Ti islands by hand and measuring their areas using imageJ [10]. Fig. S2 and S3 shows examples of these measurements.
VII.5 Raman spectroscopy
Raman spectroscopy was performed using a Renishaw inVia confocal Raman microscope with a 532 nm excitation laser. We acquired maps with 0.5 µm spacing between spectra. We then sum 10 spectra to obtain the data shown in Fig. 5.
VII.6 Atomic force microscopy
AFM was performed on a Cypher VRS AFM operated in tapping mode.
VII.7 Winterbottom reconstruction calculations
We calculated the equilibrium Winterbottom shape of Ti nanoislands on Gr using a Mathematica interface [40, 58, 69]. We obtained the relative surface energies of the crystal facets for HCP Ti from high-throughput density functional theory [70, 71, 72, 73]. The Winterbottom reconstruction shown in 3(b) was obtained using the following interface energies (all given in J/m2): $\{0001\}$: 2.15; $\{11\bar{2}1\}$: 1.93; $\{2\bar{1}\bar{1}2\}$: 2.01; $\{11\bar{2}0\}$: 2.04. A more quantitative value for the interface energy could not be obtained, since the Ti islands do not form their full equilibrium geometries, however the overall shape of the islands tends closer towards equilibrium upon heated deposition.
VII.8 Density Functional Theory Calculations
The DFT calculations are carried out using VASP 5.2 [74] employing the Perdew, Burke, and Ernzerhof (PBE) Generalized Gradient Approximation (GGA) [75] for the exchange-correlation function and PAW pseudopotentials [76, 77]. Van der Waals interactions are accounted for using the Tkatchenko-Scheffler correction [78, 79]. The electronic wavefunctions are expanded with planewaves with energy less than 400 eV. The electronic self-consistent loop is terminated when energy changes are less than $1\cdot 10^{-6}$ eV and the ionic relaxations are considered converged when the magnitude of the largest force on any atom is less than 0.01 eV/Å. The Gr slab is modeled using a 3$\times$3 surface supercell created from graphite with hexagonal lattice with lattice constants 2.458 and 6.592 Å. To mitigate the fictitious interactions between the images in the non-periodic direction perpendicular to the Gr slab, we used a supercell approach with at least 12 Å of vacuum. The Brillouin zone is sampled using the $\times 5\times$1 Monkhorst-Pack grid. All of the calculations are spin-polarized. To determine the interface, we used the Zur and McGill [80] lattice matching scheme in conjunction with lattice shifs [81, 82]. The lattice matching scheme accounts for possible rotations of one lattice with respect to the other, or expansion of both lattices, in addition to lattice reduction scheme, to minimize the strain and lattice mismatch between the two lattices. The precision of the match gives an upper bound on the strain required to form chemical bonds across the interface. The lattice shifts ensure that the atom registry at the interface is optimum [81].
VIII Acknowledgements
This work is primarily supported through the U.S. Department of Energy, Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division under FWP ERKCK47 ‘Understanding and Controlling Entangled and Correlated Quantum States in Confined Solid-state Systems Created via Atomic Scale Manipulation’, as well as partially supported by the Quantum Science Center (QSC), a National Quantum Information Research Center of the U.S. Department of Energy (DOE) on probing quantum matter. W.A.S. is partially supported by the US National Science Foundation (Award No. CBET-2130804). Computational support was provided in part by the University of Pittsburgh Center for Research Computing through the resources provided on the H2P cluster, which is supported by NSF (Award No. OAC-2117681). J.D.T. acknowledges support from Independent Research Fund Denmark through grant no. 9035-00006B. K.R. acknowledges funding from the Hugh Hampton Young Memorial Fund at MIT. J.J.P. acknowledges funding from the Natural Sciences and Engineering Research Council of Canada and MIT School of Engineering. The authors would like to acknowledge Michael Tarkanian for help in manufacturing TEM sample holders and Prof. Jeehwan Kim for equipment access. This work made use of the Materials Research Laboratory Shared Experimental Facilities at MIT, supported in part by the MRSEC Program of the National Science Foundation under Award No. DMR1419807. This work was carried out in part using MIT.nano’s facilities. This work was performed in part at the Center for Nanoscale Systems (CNS), a member of the National Nanotechnology Coordinated Infrastructure Network (NNCI), which was supported by the National Science Foundation under NSF Award No. 1541959. Figure 1(d) and 2(c) were created using Vesta [83].